# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ Abstract.
It is well known that an element of the algebra of noncommutative ***-polynomials is positive in all ***-representations if and only if it is a sum of squares. This provides an effective way to determine if a given ***-polynomial is positive, by searching through sums of squares decompositions. We show that no such procedure exists for the tensor product of two noncommutative ***-polynomial algebras: determining whether a ***-polynomial of such an algebra is positive is coRE-hard. We also show that it is coRE-hard to determine whether a noncommutative ***-polynomial is trace-positive. Our results hold if noncommutative ***-polynomial algebras are replaced by other sufficiently free algebras such as group algebras of free groups or free products of cyclic groups.
@@ 1. Introduction
Recall that a self-adjoint bounded operator $T$ on a Hilbert space $H$ is positive if $\langle v,Tv\rangle\geq 0$ for all $v\in H$. A self-adjoint element $\alpha$ of a complex ***-algebra $\mathcal{A}$ is said to be positive if $\pi(\alpha)$ is positive for all ***-representations $\pi:\mathcal{A}\to\mathcal{B}(H)$. Suppose that $\mathcal{A}$ is the complexification of a finitely-generated ***-algebra $\mathcal{A}_{\mathbb{Q}}$ over $\mathbb{Q}$. Elements of $\mathcal{A}_{\mathbb{Q}}$ can be represented as noncommutative ***-polynomials with rational coefficients, and hence we can work with them concretely, for instance on a computer.
In many situations, we’d like to be able to decide whether an element $a\in\mathcal{A}_{\mathbb{Q}}$ is positive in $\mathcal{A}$, and if it is, find a certificate of positivity.
A prototypical example is $\mathcal{A}=\mathbb{C}\mathcal{F}_{n}$, the group algebra of the free group $\mathcal{F}_{n}$, with $\mathcal{A}_{\mathbb{Q}}=\mathbb{Q}\mathcal{F}_{n}$. In this case, the following two properties hold:
@@
Property (a) is called the SOS property; for archimedean algebras (see the next section for a definition) such as $\mathbb{C}\mathcal{F}_{n}$, this property is also called archimedean closed. Since a sum of squares is always positive, a sum of squares decomposition for an element $a$ can be used as a certificate of positivity. The SOS property states that every positive element has such a certificate. For archimedean algebras, property (b) states that the enveloping $C^{*}$-algebra of $\mathcal{A}$ is residually finite-dimensional (RFD), or equivalently that the quadratic module generated by hermitian squares in $\mathcal{A}$ is RFD [ANT19]. For convenience, we’ll shorten this and just say that $\mathcal{A}$ is RFD. [^1] But note that this is not equivalent to the standard definition of RFD for ***-algebras, which says that all non-zero elements are non-zero in some finite-dimensional representation. If $\mathcal{A}$ is RFD, then for every non-positive self-adjoint element $\alpha\in\mathcal{A}$, there is a finite-dimensional representation $\phi$ such that $\phi(\alpha)$ is not positive. Thus if $\mathcal{A}_{\mathbb{Q}}$ is finitely-presented, then this representation can be used as a certificate of non-positivity. If $\mathcal{A}$ has the SOS property and is RFD, and $\mathcal{A}_{\mathbb{Q}}$ is finitely-presented, it is possible to decide whether a self-adjoint element $\alpha\in\mathcal{A}_{\mathbb{Q}}$ is positive by simultaneously searching for a sum of squares decomposition for $\alpha$, and for a finite-dimensional representation showing that $\alpha$ is not positive.
@@
For the free group algebra, property (a) is sometimes attributed to an unpublished result of Schmüdgen from the 1980s (see [NT13, Oza13]), with other proofs given in [McC01, HMP04, BT07, NT13, Oza13]. Property (b) is a result of Choi [Cho80]. Both properties also hold for the algebra of noncommutative ***-polynomials *$\mathbb{C}^{*}\langle x_{1},\ldots,x_{n}\rangle$ by theorems of Helton [Hel02] and McCullough [McC01], for the ***-algebra of $n$ contractions by theorems of Helton-McCullough-Putinar [HMP04] or Helton-Klep-McCullough [HKM12], and for the group algebra $\mathbb{C}\mathbb{Z}_{m}^{*n}$ of a free product of cyclic groups again by [HMP04] or [HKM12]. That all these algebras are RFD also follows from [GM90] or [EL92]. Thus it is possible to decide positivity for all these algebras using the algorithm above, although actually we can do better: the proofs of the SOS property above are effective, meaning that they provide bounds on the degree of polynomials in the sum of squares needed to show that an element is positive. Hence we can decide positivity by searching just through SOS decompositions and stopping when the bound is reached.
@@
In comparison, the commutative polynomial rings $\mathbb{C}[x_{1},\ldots,x_{n}]$ famously do not have the SOS property for $n\geq 2$. However, Artin’s solution of Hilbert’s 17th problem states that every positive polynomial is a sum of squares of rational functions [Art27]. More generally, a result of Scheiderer [Sch00] states that finitely generated connected commutative ***-algebras of dimension $\geq 3$, such as the group algebra $\mathbb{C}\mathbb{Z}^{n}$ for $n\geq 3$, do not have the SOS property (provided they have a state). Certificates of positivity are provided for all finitely-generated commutative algebras by the Krivine-Stengle positivstellensatz [Kri64, Ste74]. Effective versions of Artin’s theorem and the Krivine-Stengle positivstellensatz are proven in [LPR20], and although the degree bounds are too large to yield practical algorithms, this gives a way to decide positivity in principle. For finitely-generated commutative algebras, positivity can also be decided with Tarksi quantifier elimination or other algorithms for the first-order theory of the reals. In fact, since non-positivity lies in the existential theory of the reals, positivity can be decided in PSPACE [Can88].
@@
If $\mathcal{A}$ is a finitely-generated commutative archimedean ***-algebra, then Putinar’s positivstellensatz states that every element which is strictly positive on the state space of $\mathcal{A}$ is a sum of squares [Put93]. For all archimedean ***-algebras $\mathcal{A}$, Helton and McCullough’s positivstellensatz [HM04] says that $\alpha\in\mathcal{A}$ is positive if and only if $\alpha+\epsilon$ is a sum of squares for all $\epsilon>0$ (this exact formulation seems to appear first in [Cim09], see also [Sch09, Oza13]). This does not give a method for deciding positivity, but does show that deciding positivity is in the class $\Pi_{2}^{0}$ for archimedean algebras.
@@
While positivity is decidable for the free group algebras and other prototypical noncommutative ***-algebras $\mathcal{A}$ listed above, the point of this paper is to show that positivity becomes undecidable for tensor products $\mathcal{A}\otimes\mathcal{A}$ of such algebras:
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}=\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathbb{C}$ be the complexification. Then the decision problem
@@
Specifically, we show that there is a computable mapping from Turing machines $M$ to self-adjoint elements $\alpha_{M}\in\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$, such that $\alpha_{M}$ is positive if and only if $M$ does not halt. The bounds on $n$ and $m$ in Theorem 1.1 are tight, as the algebras *$\mathbb{Q}^{*}\langle x\rangle\otimes\mathbb{Q}^{*}\langle x\rangle=\mathbb{Q}[x_{1},x_{2}]$, $\mathbb{Q}\mathcal{F}_{1}\otimes\mathbb{Q}\mathcal{F}_{1}=\mathbb{Q}\mathbb{Z}\times\mathbb{Z}$, and $\mathbb{Q}\mathbb{Z}_{m}\times\mathbb{Q}\mathbb{Z}_{m}$ are commutative, and positivity is also decidable for $\mathbb{Q}\mathbb{Z}_{2}^{*2}\times\mathbb{Z}_{2}^{*2}$ (see Example 6.8). As an immediate corollary of Theorem 1.1, we get:
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}$ be the complexification. Then there are elements of $\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$ which are positive, but are not sums of hermitian squares in $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$, so $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$ does not have the SOS property.
@@
The proof of Theorem 6.1 is by reduction to tracial positivity. An element $\alpha$ of a ***-algebra $\mathcal{A}$ is said to be trace-positive if $\tau(\alpha)\geq 0$ for all (bounded) tracial states $\tau$ on $\mathcal{A}$. We show that tracial positivity is undecidable for free ***-polynomials and group algebras:
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{m}^{*N}$ for $m\geq 2$. If $N$ is large enough, then the decision problem
@@
Klep and Schweighofer [KS08] prove an analog of Helton and McCullough’s positivstellensatz for the algebra of bounded contractions: an element $\alpha$ is trace-positive if and only if $\alpha+\epsilon$ is a sum of hermitian squares and commutators for all $\epsilon>0$. This is extended to free group algebras by Juschenko and Popovych [JP11], to archimedean algebras by Ozawa [Oza13], and to *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ by Klep, Scheiderer, and Volčič [KSV23]. Quarez [Qua15] (see also [KSV23, Example 3.2]) gives examples of trace-positive polynomials in *$\mathbb{Q}^{*}\langle x_{1},x_{2}\rangle$ which are not sums of hermitian squares and commutators, but which are sums of squares and commutators of noncommutative rational functions. As in 1.2, Theorem 1.3 immediately implies that all the algebras listed in the theorem contain trace-positive elements which are not sums of squares and commutators of noncommutative rational functions. Unlike Theorem 1.1, we do not know the smallest bound $N$ such that Theorem 1.3 holds.
@@
Part of the motivation for Theorem 1.1 comes from Kirchberg’s reformulation of the Connes embedding problem, sometimes called Kirchberg’s conjecture [Kir93], and the recent MIP*=RE theorem of Ji, Natarajan, Vidick, Wright, and Yuen [${}^{+}$20], which resolves this conjecture. Kirchberg’s conjecture asks whether the max and min tensor norms on $\mathbb{C}\mathcal{F}_{n}\otimes\mathcal{F}_{n}=\mathbb{C}\mathcal{F}_{n}\times\mathcal{F}_{n}$ agree for $n\geq 2$ [Kir93]. As Ozawa observes [Oza04, Oza13], if $\mathcal{A}$ is archimedean and RFD, then the max and min tensor norms on $\mathcal{A}\otimes\mathcal{A}$ agree if and only if $\mathcal{A}\otimes\mathcal{A}$ is RFD, so Kirchberg’s conjecture is equivalent to whether $\mathbb{C}\mathcal{F}_{n}\otimes\mathbb{C}\mathcal{F}_{n}=\mathbb{C}\mathcal{F}_{n}\times\mathcal{F}_{n}$ is RFD. Kirchberg’s conjecture is still equivalent to the Connes embedding problem if $\mathcal{F}_{n}$ is replaced by $\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$ [Oza13, Fri12]. The algebras $\mathbb{C}\mathbb{Z}_{m}^{*n}\otimes\mathbb{C}\mathbb{Z}_{m}^{*n}$ are of particular interest in quantum information, since they correspond to bipartite (meaning composed of two spatially separated subsystems) measurement scenarios with $n$ measurement settings and $m$ measurement outcomes per subsystem, and they are closely connected with the complexity class MIP* of multiprover interactive proof systems with entangled provers.
@@
The main result of [${}^{+}$20] is that MIP* is equal to RE, the class of recursively enumerable languages. In algebraic language, this implies that there is a computable mapping from Turing machines $M$ to elements $\alpha_{M}\in\overline{\mathbb{Q}}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ (where $\overline{\mathbb{Q}}$ is the algebraic closure of $\mathbb{Q}$, and $m$ and $n$ also depend on $M$), such that if $M$ halts then $\alpha_{M}$ is not positive in finite-dimensional representations of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$, and if $M$ does not halt, then $\alpha_{M}-1$ is positive in finite-dimensional representations of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$. This shows that the promise problem
@@
Note that $\alpha-1$ is positive if and only if $\alpha$ is strictly positive with spectrum contained in $[1,+\infty)$, so (GapPos) is a “gapped” version of the decision problem (Pos) considered in Theorem 1.1 (albeit with $\mathbb{Q}$ replaced by $\overline{\mathbb{Q}}$, and $n$ and $m$ allowed to vary). By Helton and McCullough’s positivstellensatz, if $\alpha-1$ is positive then $\alpha$ has an SOS decomposition proving positivity, so unlike (Pos), the problem (GapPos) is contained in RE. Since RE does not contain coRE, (GapPos) cannot be coRE-hard, and hence the MIP*=RE theorem implies that $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ is not RFD for large enough $n$ and $m$.
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}$ be the complexification. Then the decision problem (Pos) from Theorem 1.1 is $\Pi_{2}^{0}$-complete.
Tracial positivity has also been studied in connection with the Connes embedding problem [Had01, KS08, BDKS14, JP11, Oza13], and the above connection with $\text{MIP}^{*}=\text{RE}$ and $\text{MIP}^{\text{co}}=\text{coRE}$ could be described equally as well in terms of tracial positivity and MIP protocols with synchronous strategies (see e.g. [HMPS19, KPS18, MNY22, NZ23]). For tracial positivity, we make the following conjecture.
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras *$\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{m}^{*N}$ for $m\geq 2$, and let $\mathcal{A}$ be the complexification. If $N$ is large enough, then the decision problem (TrPos) from Theorem 1.3 is $\Pi_{2}^{0}$-complete.

### 18 Mar 2026 v2

@@ Abstract.
It is well known that an element of the algebra of noncommutative *-polynomials is positive in all *-representations if and only if it is a sum of squares. This provides an effective way to determine if a given *-polynomial is positive, by searching through sums of squares decompositions. We show that no such procedure exists for the tensor product of two noncommutative *-polynomial algebras: determining whether a *-polynomial of such an algebra is positive is coRE-hard. We also show that it is coRE-hard to determine whether a noncommutative *-polynomial is trace-positive. Our results hold if noncommutative *-polynomial algebras are replaced by other sufficiently free algebras such as group algebras of free groups or free products of cyclic groups.
@@ 1. Introduction
Recall that a self-adjoint bounded operator $T$ on a Hilbert space $H$ is positive if $\langle v,Tv\rangle\geq 0$ for all $v\in H$. A self-adjoint element $\alpha$ of a complex *-algebra $\mathcal{A}$ is said to be positive if $\pi(\alpha)$ is positive for all *-representations $\pi:\mathcal{A}\to\mathcal{B}(H)$. Suppose that $\mathcal{A}$ is the complexification of a finitely generated *-algebra $\mathcal{A}_{\mathbb{Q}}$ over $\mathbb{Q}$. Elements of $\mathcal{A}_{\mathbb{Q}}$ can be represented as noncommutative *-polynomials with rational coefficients, and hence we can work with them concretely, for instance on a computer.
In many situations, we’d like to be able to decide whether an element $a\in\mathcal{A}_{\mathbb{Q}}$ is positive in $\mathcal{A}$, and if it is, find a certificate of positivity. A prototypical example is $\mathcal{A}=\mathbb{C}\mathcal{F}_{n}$, the group algebra of the free group $\mathcal{F}_{n}$, with $\mathcal{A}_{\mathbb{Q}}=\mathbb{Q}\mathcal{F}_{n}$. In this case, the following two properties hold:
@@
Property (a) is called the SOS property; for archimedean algebras (see the next section for a definition) such as $\mathbb{C}\mathcal{F}_{n}$, this property is also called archimedean closed. Since a sum of squares is always positive, a sum of squares decomposition for an element $a$ can be used as a certificate of positivity. The SOS property states that every positive element has such a certificate. For archimedean algebras, property (b) states that the enveloping $C^{*}$-algebra of $\mathcal{A}$ is residually finite-dimensional (RFD), or equivalently that the quadratic module generated by hermitian squares in $\mathcal{A}$ is RFD [ANT19]. For convenience, we’ll shorten this and just say that $\mathcal{A}$ is RFD. [^1] But note that this is not equivalent to the standard definition of RFD for *-algebras, which says that all non-zero elements are non-zero in some finite-dimensional representation. If $\mathcal{A}$ is RFD, then for every non-positive self-adjoint element $\alpha\in\mathcal{A}$, there is a finite-dimensional representation $\phi$ such that $\phi(\alpha)$ is not positive. Thus if $\mathcal{A}_{\mathbb{Q}}$ is finitely presented, then this representation can be used as a certificate of non-positivity. If $\mathcal{A}$ has the SOS property and is RFD, and $\mathcal{A}_{\mathbb{Q}}$ is finitely presented, it is possible to decide whether a self-adjoint element $\alpha\in\mathcal{A}_{\mathbb{Q}}$ is positive by simultaneously searching for a sum of squares decomposition for $\alpha$, and for a finite-dimensional representation showing that $\alpha$ is not positive.
@@
For the free group algebra, property (a) is sometimes attributed to an unpublished result of Schmüdgen from the 1980s (see [NT13, OZA13]), and can also be found in [MCC01, HMP04, BT07, NT13, OZA13]. Property (b) is a result of Choi [CHO80]. Both properties also hold for the *-algebra of noncommutative *-polynomials $\mathbb{C}^{*}\langle x_{1},\ldots,x_{n}\rangle$ by theorems of Helton [HEL02] and McCullough [MCC01], and for the *-algebra of $n$ contractions by theorems of Helton-McCullough-Putinar [HMP04] or Helton-Klep-McCullough [HKM12]. For the group algebra of a virtually free group such as the free product $\mathbb{Z}_{m}^{*n}$ of cyclic groups, a proof of both properties is again attributed to Schmüdgen [NT13], with a complete proof appearing in [COH23]. That all these algebras are RFD also follows from [GM90] or [EL92]. Thus it is possible to decide positivity for all these algebras using the algorithm above, although actually we can do better: the proofs of the SOS property above are effective, meaning that they provide bounds on the degree of polynomials in the sum of squares needed to show that an element is positive. Hence we can decide positivity by searching just through SOS decompositions and stopping when the bound is reached.
@@
In comparison, the commutative polynomial rings $\mathbb{C}[x_{1},\ldots,x_{n}]$ famously do not have the SOS property for $n\geq 2$. However, Artin’s solution of Hilbert’s 17th problem states that every positive polynomial is a sum of squares of rational functions [ART27]. More generally, a result of Scheiderer [SCH00] states that finitely generated connected commutative *-algebras of dimension $\geq 3$, such as the group algebra $\mathbb{C}\mathbb{Z}^{n}$ for $n\geq 3$, do not have the SOS property (provided they have a state). Certificates of positivity are provided for all finitely generated commutative algebras by the Krivine-Stengle Positivstellensatz [KRI64, STE74]. Effective versions of Artin’s theorem and the Krivine-Stengle Positivstellensatz are proven in [LPR20], and although the degree bounds are too large to yield practical algorithms, this gives a way to decide positivity in principle. For finitely generated commutative algebras, positivity can also be decided with Tarski quantifier elimination or other algorithms for the first-order theory of the reals. In fact, since non-positivity lies in the existential theory of the reals, positivity can be decided in PSPACE [CAN88].
@@
If $\mathcal{A}$ is a finitely generated commutative archimedean *-algebra, then Schmüdgen’s Positivstellensatz states that every element which is strictly positive on the state space of $\mathcal{A}$ is a sum of squares [SCH91]. For all archimedean *-algebras $\mathcal{A}$, Helton and McCullough’s Positivstellensatz [HM04] says that $\alpha\in\mathcal{A}$ is positive if and only if $\alpha+\epsilon$ is a sum of squares for all $\epsilon>0$ (this exact formulation seems to appear first in [CIM09], see also [SCH09, OZA13]). This does not give a method for deciding positivity, but does show that deciding positivity is in the class $\Pi_{2}^{0}$ for finitely presented archimedean algebras.
@@
While positivity is decidable for the free group algebras and other prototypical noncommutative *-algebras $\mathcal{A}$ listed above, the point of this paper is to show that positivity becomes undecidable for tensor products $\mathcal{A}\otimes\mathcal{A}$ of such algebras:
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}=\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathbb{C}$ be the complexification. Then the decision problem
@@
Here, $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ denotes the *-algebra of noncommutative *-polynomials over $\mathbb{Q}$. Specifically, Theorem 1.1 shows that there is a computable mapping from Turing machines $M$ to self-adjoint elements $\alpha_{M}\in\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$, such that $\alpha_{M}$ is positive if and only if $M$ does not halt. The bounds on $n$ and $m$ in Theorem 1.1 are tight, as the algebras $\mathbb{Q}^{*}\langle x\rangle\otimes\mathbb{Q}^{*}\langle x\rangle=\mathbb{Q}[x_{1},x_{2}]$, $\mathbb{Q}\mathcal{F}_{1}\otimes\mathbb{Q}\mathcal{F}_{1}=\mathbb{Q}\mathbb{Z}\times\mathbb{Z}$, and $\mathbb{Q}\mathbb{Z}_{m}\times\mathbb{Q}\mathbb{Z}_{m}$ are commutative, and positivity is also decidable for $\mathbb{Q}\mathbb{Z}_{2}^{*2}\times\mathbb{Z}_{2}^{*2}$ (see Example 6.8). As an immediate corollary of Theorem 1.1, we get:
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}$ be the complexification. Then there are elements of $\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$ which are positive, but are not sums of hermitian squares in $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$, so $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$ does not have the SOS property.
@@
The proof of Theorem 6.1 is by reduction to tracial positivity. An element $\alpha$ of a *-algebra $\mathcal{A}$ is said to be trace-positive if $\tau(\alpha)\geq 0$ for all (bounded) tracial states $\tau$ on $\mathcal{A}$. We show that tracial positivity is undecidable for free *-polynomials and group algebras:
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{m}^{*N}$ for $m\geq 2$. If $N$ is large enough, then the decision problem
@@
Klep and Schweighofer [KS08] prove an analog of Helton and McCullough’s Positivstellensatz for the algebra of bounded contractions: an element $\alpha$ is trace-positive if and only if $\alpha+\epsilon$ is a sum of hermitian squares and commutators for all $\epsilon>0$. This is extended to free group algebras by Juschenko and Popovych [JP11], to archimedean algebras by Ozawa [OZA13], and to $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ by Klep, Scheiderer, and Volčič [KSV23]. Quarez [QUA15] (see also [KSV23, Example 3.2]) gives examples of trace-positive polynomials in $\mathbb{Q}^{*}\langle x_{1},x_{2}\rangle$ which are not sums of hermitian squares and commutators, but which are sums of squares and commutators of noncommutative rational functions. As in 1.2, Theorem 1.3 immediately implies that all the algebras listed in the theorem contain trace-positive elements which are not sums of squares and commutators of noncommutative rational functions. Unlike Theorem 1.1, we do not know the smallest bound $N$ such that Theorem 1.3 holds.
@@
Part of the motivation for Theorem 1.1 comes from Kirchberg’s reformulation of the Connes embedding problem, sometimes called Kirchberg’s conjecture [KIR93], and the recent MIP*=RE theorem of Ji, Natarajan, Vidick, Wright, and Yuen [JNV+20], which resolves this conjecture. Kirchberg’s conjecture asks whether the max and min tensor norms on $\mathbb{C}\mathcal{F}_{n}\otimes\mathcal{F}_{n}=\mathbb{C}\mathcal{F}_{n}\times\mathcal{F}_{n}$ agree for $n\geq 2$ [KIR93]. As Ozawa observes [OZA04, OZA13], if $\mathcal{A}$ is archimedean and RFD, then the max and min tensor norms on $\mathcal{A}\otimes\mathcal{A}$ agree if and only if $\mathcal{A}\otimes\mathcal{A}$ is RFD, so Kirchberg’s conjecture is equivalent to whether $\mathbb{C}\mathcal{F}_{n}\otimes\mathbb{C}\mathcal{F}_{n}=\mathbb{C}\mathcal{F}_{n}\times\mathcal{F}_{n}$ is RFD. Kirchberg’s conjecture is still equivalent to the Connes embedding problem if $\mathcal{F}_{n}$ is replaced by $\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$ [OZA13, FRI12]. The algebras $\mathbb{C}\mathbb{Z}_{m}^{*n}\otimes\mathbb{C}\mathbb{Z}_{m}^{*n}$ are of particular interest in quantum information, since they correspond to bipartite (meaning composed of two spatially separated subsystems) measurement scenarios with $n$ measurement settings and $m$ measurement outcomes per subsystem, and they are closely connected with the complexity class MIP* of multiprover interactive proof systems with entangled provers.
@@
The main result of [JNV+20] is that MIP* is equal to RE, the class of recursively enumerable languages. In algebraic language, this implies that there is a computable mapping from Turing machines $M$ to elements $\alpha_{M}\in\overline{\mathbb{Q}}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ (where $\overline{\mathbb{Q}}$ is the algebraic closure of $\mathbb{Q}$, and $m$ and $n$ also depend on $M$), such that if $M$ halts then $\alpha_{M}$ is not positive in finite-dimensional representations of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$, and if $M$ does not halt, then $\alpha_{M}-1$ is positive in finite-dimensional representations of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$. This shows that the promise problem
@@
Note that $\alpha-1$ is positive if and only if $\alpha$ is strictly positive with spectrum contained in $1,+\infty)$, so (GapPos) is a “gapped” version of the decision problem (Pos) considered in Theorem 1.1 (albeit with $\mathbb{Q}$ replaced by $\overline{\mathbb{Q}}$, and $n$ and $m$ allowed to vary). By Helton and McCullough’s Positivstellensatz, if $\alpha-1$ is positive then $\alpha$ has an SOS decomposition proving positivity. Due to the promise in (GapPos), a proof that $\alpha$ is positive is also a proof that $\alpha-1$ is positive. So unlike (Pos), the problem (GapPos) is contained in RE. Since RE does not contain coRE, (GapPos) cannot be coRE-hard, and hence the MIP*=RE theorem implies that $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ is not RFD for large enough $n$ and $m$.
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}$ be the complexification. Then the decision problem (Pos) from Theorem 1.1 is $\Pi_{2}^{0}$-complete.
Tracial positivity has also been studied in connection with the Connes embedding problem [HAD01, KS08, BDK+14, JP11, OZA13], and the above connection with $\text{MIP}^{*}=\text{RE}$ and $\text{MIP}^{\text{co}}=\text{coRE}$ could be described equally as well in terms of tracial positivity and MIP protocols with synchronous strategies (see e.g. [HMP+19, KPS18, MNY22, NZ23]). For tracial positivity, we make the following conjecture.
@@
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{m}^{*N}$ for $m\geq 2$, and let $\mathcal{A}$ be the complexification. If $N$ is large enough, then the decision problem (TrPos) from Theorem 1.3 is $\Pi_{2}^{0}$-complete.